OpenStudy (anonymous):

The minimum value of (sinx+cosecx)^2 + (cosx+secx)^2

mathslover (mathslover):

@yrajyalakshmi - In order to find the minimum value of \((\sin x + \csc x)^2 + (\cos x + \sec x) ^2 \) You should first expand the terms using \((a+b)^2 = a^2+b^2+2ab\) \(\sin^2 x + \csc ^2 x + 2\sin x \csc x + \cos^2x + \sec^2 x +2\cos x \sec x\) (mark this as equation 1) Now, since, \(\sin x = \cfrac{1}{\csc x} \\ \text{OR}\\ \sin x \times \csc x = 1 \) Similarly, \(\cos x = \cfrac{1}{\sec x} \\ \text{OR} \\ \cos x \sec x = 1\) Therefore, equation one becomes : \(\sin^2 (x) + \csc^2 (x) + 2(1) + \cos^2 (x) + \sec^2 (x) + 2(1) \\ \text{Rearrange this accordingly} \\ \sin^2(x) + \cos^2(x) + \csc^2(x) + \sec^2 (x) + 4 \) You should remember the identity that : \(\sin^2 (x) + \cos^2 (x) = 1\) Therefore, we get : \(1 + 4 + \csc ^2(x) + \sec^2(x) \) \(5 + \csc^2(x) + \sec^2(x) \) Now, I can write \(\csc^2 (x) \) as \(\cfrac{1}{\sin^2 (x)}\) and \(\sec^2(x) \) as\(\cfrac{1}{\cos^2(x)}\) Therefore, you get : \(5 + \cfrac{1}{\sin^2(x)} + \cfrac{1}{\cos^2(x)} \\ 5 + \cfrac{\cos^2(x) + \sin^2(x) }{\cos^2(x) \times \sin^2(x)}\) Again \(\cos^2(x) + \sin^2(x) = 1\) Therefore, you get : \(5 + \cfrac{1}{\cos^2(x) \times \sin^2(x)}\) Now, we will be using sine product formula : \(\sin (2x) = 2\sin x \cos x\\ \text{OR} \\ \cfrac{1}{2} \sin(2x) = \sin x \cos x\) Therefore, \(5 + \cfrac{1}{\cos^2(x) \times \sin^2(x)}\) can be written as : \(5 + \cfrac{1}{ (\cos (x) \times \sin(x) )^2 } = 5 + \cfrac{1}{(\frac{1}{2} \sin (2x) )^2}\) \( 5 + \cfrac{4}{(\sin(2x))^2}\) Minimum value of sin(2x) = 1 ( in case of zero, it will be undefined , so, we will not consider the case of x = 0 .) Put that minimum value at the place of \((\sin(2x))^2\) and thus, you will get the minimum value of the given expression.

mathslover (mathslover):

While , there is a much easier method for this, called AM-GM Inequality. Have you studied it? I would be able to explain you only when you're available online. Good Luck!